OpenStudy (anonymous):

can someone guide me the right way please? use set-builder notation to write the solution set 3y + x = 4 x = 2y − 1

OpenStudy (kinggeorge):

Do you know what set-builder notation?

OpenStudy (anonymous):

Zed if i can get your assistance real quick it would be helpful

OpenStudy (anonymous):

if it has an infinite number of solutions

OpenStudy (anonymous):

is that what you meant King?

OpenStudy (anonymous):

or if I know what set builder notation is?

OpenStudy (kinggeorge):

what the notation actually is.

OpenStudy (anonymous):

no it doesn't say

OpenStudy (anonymous):

do I just plug in the 2nd equation into the first?

OpenStudy (kinggeorge):

have you seen solutions posted in a similar fashion to these?\[\left\{ x, y \in R: x, y > 0 \right\}\]\[\left\{x, y \in R: x \geq 5, y < 0 \right\}\]

OpenStudy (kinggeorge):

That's what set-builder notation looks like.

OpenStudy (kinggeorge):

As for solving the equation itself, plugging the 2nd equation into the first should work.